# Stability of flocking in the reciprocal two-species Vicsek model: Effects of relative population, motility, and noise

### v1

@@ IV.4 Noise heterogeneity, or unfriendly ?hot? and ?cold? particles
?figurename? 18: (color online) Phase diagrams for noise heterogeneity. (a) $\eta-\Delta\eta$ phase diagram for fixed $\rho=1.5$; (b) $\rho-\Delta\eta$ phase diagram for fixed $\eta=0.3$. For both cases, the velocity moduli is $v_{0}=0.5$. The boundary lines act as a guide to the eyes.
@@ V Discussion
For strong motility heterogeneity, we established the absence of the PF state in the coexistence regime as the spatial segregation between species is compromised. However, considering activity landscapes with region-dependent motilities, the dynamical behavior contrasts with the simple motility heterogeneity. We find high retention of the PF behavior for a given geometry, where the fast region is larger than the slow region, and the emergence of a vertical APF state for large relative velocity, due to particle trapping in the slow region. This shows that interruption by environmental constructs plays a big role in shaping the nature of flocking.
We also find that species motility is significantly affected by noise heterogeneity, where the colder species (subjected to lower noise) moves faster than the hotter one due to its higher band velocity, eventually catching up and transiently disrupting any PF structure. However, the pattern re-emerges, reflecting a dynamic yet robust response to noise asymmetry. This behavior contrasts motility heterogeneity, where differing species velocities lead to the absence of the PF state in the coexistence regime.
@@
In confined environments, stability emerges more easily if components interact frequently, but it can also be disrupted by the surrounding ?habitat?. For instance, if one species relies on a resource for growth while the other does not, this imbalance can induce indirect inter-species interactions. A classic case is antagonistic predator-prey dynamics: if a predator consumes prey that depends on an external resource (e.g., vegetation or water), fluctuations in that resource indirectly affect the predator, even though the predator interacts only with the prey. Our framework could be extended to capture such environmental feedback by introducing a localized, depletable resource field that influences only one species. Furthermore, non-uniform system noise and obstacles Martinez et al. (2018) can be considered as other convincing candidates for imparting spatial heterogeneity. Another natural extension would be to let particle velocities depend on the local crowding by introducing a density-dependent self-propulsion speed similar to the softcore restriction considered in Ref. Karmakar et al. (2023). Finally, the TSVM Chatterjee et al. (2023) does not have the factor of agent size. We can bring size (and later, even shape) into the picture by first working with finite-size hard discs Martinez et al. (2018) instead of point particles in the future.
@@ Acknowledgements
  • Mangeat et al. (2024) M. Mangeat, S. Chatterjee, J. D. Noh,  and H. Rieger, ?Emergent complex phases in a discrete flocking model with reciprocal and non-reciprocal interactions,? arXiv preprint arXiv:2412.02501  (2024).
@@
  • Dutta et al. (2025) A. K. Dutta, M. Mangeat, H. Rieger, R. Paul, and S. Chatterjee,  ?Supplementary movies for “The heterogeneous two-species Vicsek model”,? https://doi.org/10.5281/zenodo.15241518, Zenodo  (2025).
  • (53) See Supplementary Material.
@@
Supplemental Material for “Stability of flocking in the reciprocal two-species Vicsek model: Effects of relative population, motility, and noise”
Aditya Kumar Dutta,1 Matthieu Mangeat,2 Heiko Rieger,2 Raja Paul,1 and Swarnajit Chatterjee2
1School of Mathematical & Computational Sciences, Indian Association for the Cultivation of Science, Kolkata – 700032, India.
[…]
Description of Movies
  1. 1.
    Movie S1: Movie (movieS1_PHTSVM_proba.mp4) corresponds to Fig. 3 showing the probability distribution $P(v_{s},v_{a})$ for varying population heterogeneity $m_{0}$. Parameters: $\rho=0.5$, $\eta=0.24$, $v_{0}=0.5$, $L_{x}=256$, and $L_{y}=32$.
  2. 2.
    Movie S2: Movie (movieS2_MHTSVM_proba.mp4) corresponds to Fig. 8 showing the probability distribution $P(v_{s},v_{a})$ for varying motility heterogeneity $\Delta v$. Parameters: $\rho=0.5$, $\eta=0.24$, $v_{0}=0.5$, $L_{x}=256$, and $L_{y}=32$.
  3. 3.
    Movie S3: Movie (movieS3_SHTSVM_proba_vR.mp4) corresponds to Fig. 13(a–e) showing the probability distribution $P(v_{s},v_{a})$ for spatial heterogeneity and varying $v_{\rm R}$. Parameters: $\rho=0.5$, $\eta=0.24$, $v_{\rm B}=0.5$, $\zeta=0.5$, $L_{x}=256$, and $L_{y}=32$.
  4. 4.
    Movie S4: Movie (movieS4_SHTSVM_proba_zeta.mp4) corresponds to Fig. 13(f–j) showing the probability distribution $P(v_{s},v_{a})$ for spatial heterogeneity and varying $\zeta$. Parameters: $\rho=0.5$, $\eta=0.24$, $v_{\rm B}=0.5$, $v_{\rm R}=0.2$, $L_{x}=256$, and $L_{y}=32$.
  5. 5.
    Movie S5: Movie (movieS5_NHTSVM_proba.mp4) corresponds to Fig. 16 showing the probability distribution $P(v_{s},v_{a})$ for varying noise heterogeneity $\Delta\eta$. Parameters: $\rho=0.5$, $\eta=0.24$, $v_{0}=0.5$, $L_{x}=256$, and $L_{y}=32$.

### v2

@@ IV.4 Noise heterogeneity, or unfriendly ?hot? and ?cold? particles
?figurename? 18: (color online) Phase diagrams for noise heterogeneity. (a) $\eta-\Delta\eta$ phase diagram for fixed $\rho=1.5$; (b) $\rho-\Delta\eta$ phase diagram for fixed $\eta=0.3$. For both cases, the velocity modulus is $v_{0}=0.5$. The boundary lines act as a guide to the eyes.
@@ V Discussion
For strong motility heterogeneity, we established the absence of the PF state in the coexistence regime as the spatial segregation between species is compromised. However, considering activity landscapes with region-dependent motilities, the dynamical behavior contrasts with the simple motility heterogeneity. We find high retention of the PF behavior for a given geometry, where the fast region is larger than the slow region, and the emergence of a vertical APF state for large relative velocity, due to particle trapping in the slow region. This shows that interruption by environmental constructs plays a big role in shaping the nature of flocking. In this regard, our current implementation is based on sharp spatial variations in motility. A natural extension would be to incorporate smooth motility gradients or time-evolving landscapes, which more accurately reflect realistic environments. Gradual changes in activity could soften inter-species collisions and mitigate sharp density mismatches, while temporally varying or fluctuating landscapes may act as continuous sources of disorder, disrupting segregation and potentially stabilizing new dynamic patterns. Exploring such scenarios presents an interesting direction for future work.
We also find that species motility is significantly affected by noise heterogeneity, where the colder species (subjected to lower noise) moves faster than the hotter one due to its higher band velocity, eventually catching up and transiently disrupting any PF structure. However, the pattern re-emerges, reflecting a dynamic yet robust response to noise asymmetry. This behavior contrasts with motility heterogeneity, where differing species velocities lead to the absence of the PF state in the coexistence regime.
@@
In confined environments, stability emerges more easily if components interact frequently, but it can also be disrupted by the surrounding ?habitat?. For instance, if one species relies on a resource for growth while the other does not, this imbalance can induce indirect inter-species interactions. A classic case is antagonistic predator-prey dynamics: if a predator consumes prey that depends on an external resource (e.g., vegetation or water), fluctuations in that resource indirectly affect the predator, even though the predator interacts only with the prey. Our framework could be extended to capture such environmental feedback by introducing a localized, depletable resource field that influences only one species. Another natural extension would be to incorporate individual-level feedback mechanisms such as quorum sensing, where particles adjust their motility or alignment in response to local density or species composition. Within the framework of our model, this can be achieved by allowing each particle’s self-propulsion speed or alignment strength to depend dynamically on local crowding or the relative concentration of different species.
Furthermore, non-uniform system noise and obstacles Martinez et al. (2018) can be considered as other convincing candidates for imparting spatial heterogeneity. Another natural extension would be to let particle velocities depend on the local crowding by introducing a density-dependent self-propulsion speed similar to the softcore restriction considered in Ref. Karmakar et al. (2023). Finally, the TSVM Chatterjee et al. (2023) does not have the factor of agent size. We can bring size (and later, even shape) into the picture by first working with finite-size hard discs Martinez et al. (2018) instead of point particles in the future.
?appendixname? A On the stability and translational order of the micro-phase separated state
?figurename? 19: Comparison of band configuration at (a) $t=10^{6}$ and (b) $t=10^{7}$ showing stable band number over long times. Parameters: $\rho=1,\eta=0.3,v_{0}=0.5$.
Here, we discuss the stability of the micro-phase separated bands and examine whether they exhibit any translational order, focusing on the band structures obtained in the presence of population heterogeneity (Sec. IV.1). We find that the micro-phase separated coexistence phase corresponds to a stable steady state, and that band coarsening is limited: the number and size of bands remain finite even at long times. Fig. 19 shows steady-state band configurations for different values of $m_{0}$ at $t=10^{6}$ and $t=10^{7}$, and confirms that once the system reaches the phase-separated state, the average number of bands remains largely unchanged over time. It is important to note that in both the VM and TSVM, fluctuations—arising from noise and finite system size—can affect band number and spacing by inducing mergers, splits, or positional shifts, leading to temporal variations in the band structure even after apparent phase separation. In a two-species system with reciprocal interactions, as in the present model, the alignment dynamics between species introduces further complexity by coupling their spatial organization.
?figurename? 20: Structure factors (non-normalized) constructed from the spatial Fourier transform of the one-dimensional projected density profile along $x$ axis: (a) $m_{0}=0.6$. (b) $m_{0}=0.8$. All prominent Bragg peaks have been marked with red dashed lines. Parameters: $\rho=1,\eta=0.3,v_{0}=0.5$.
To investigate the presence of long-range translational order similar to a flying smectic, we compute the static structure factor corresponding to the longitudinal particle density profile. This analysis is limited to cases with $m_{0}\geq 0.6$, where the minority species no longer forms distinct bands, and the overall spatial organization becomes more regular. The system is divided along the longitudinal ($x$) direction into $L_{y}$ stripes, and the one-dimensional density profile $\rho(i)$ is computed. From this, we determine the density fluctuations $\delta\rho(i)=\rho(i)-\langle\rho\rangle$, and apply a discrete Fourier transform to obtain the spectral density $\tilde{\rho}(k_{x})$. The non-normalized structure factor is then calculated as $S(k_{x})=|\tilde{\rho}(k_{x})|^{2}$. Fig. 20 presents the resulting structure factors for $m_{0}=0.6$ and $m_{0}=0.8$. In both cases, a prominent Bragg peak is observed and aligns with the expected wavevector $k_{1}=2\pi/a$, where $a=L_{x}/n_{b}$ denotes the average distance between bands and $n_{b}$ is the number of observed bands. However, no higher-order harmonics at integer multiples of $k_{1}$ are seen, and the primary peak shows considerable broadening, especially at $m_{0}=0.6$. These observations indicate notable fluctuations in both band spacing and width, inconsistent with true long-range translational order. Therefore, although the bands may appear quasi-periodic over limited regions and timescales, the system remains in a fluctuating, dynamic banded state rather than forming a true smectic.
Therefore, our analysis reveals that the micro-phase separated state in the heterogeneous TSVM exhibits finite band structures that persist over long times without coarsening into a macroscopic phase-separated state. However, the presence of broadened, irregular Bragg peaks in the structure factor signals the absence of long-range translational symmetry, distinguishing these banded states from flying smectics.
@@ Acknowledgements
  • Mangeat et al. (2025) M. Mangeat, S. Chatterjee, J. D. Noh,  and H. Rieger, ?Emergent complex phases in a discrete flocking model with reciprocal and non-reciprocal interactions,? Commun. Phys. 8, 186 (2025).
@@
  • Lardet et al. (2025) E. Lardet, L. Chen,  and T. Bertrand, ?Flocking beyond one species: Novel phase coexistence in a generalized two-species vicsek model,? arXiv preprint arXiv:2503.17617  (2025).
@@
  • Dutta et al. (2025) A. K. Dutta, M. Mangeat, H. Rieger, R. Paul,  and S. Chatterjee, ?Supplementary movies for Stability of flocking in the reciprocal two-species Vicsek model: Effects of relative population, motility, and noise,? Zenodo  (2025), 10.5281/zenodo.15241518.
  • Chen and Toner (2013) L. Chen and J. Toner, ?Universality for moving stripes: A hydrodynamic theory of polar active smectics,? Phys. Rev. Lett. 111, 088701 (2013).
  • Adhyapak et al. (2013) T. C. Adhyapak, S. Ramaswamy,  and J. Toner, ?Live soap: stability, order, and fluctuations in apolar active smectics,? Phys. Rev. Lett. 110, 118102 (2013).
  • (56) See Supplemental Material.